OpenStudy (anonymous):

8 less than the product of 5 and-4

OpenStudy (deepolisnoob):

8 less, -8, than the product of 5 and -4, 5*-4. Combine them all, youll have -4*5-8. Then solve for the problem, -4*5=-20, -20-8=-28